A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is 35% salt and Solution B is 85% salt. She wants to obtain 130 ounces of a mixture that is 80% salt. How many ounces of each solution should she use?
Found 4 solutions by ikleyn, Alan3354, greenestamps, josgarithmetic:
Answer by ikleyn(52787)

.

The maximal salt concentration in water is 28%.

https://en.wikipedia.org/wiki/Saline_water


The rest is your fantasy (or fantasy your teacher, who does not know Science).


---------------
Comment from student: The fact that the word water appears anywhere in the question is your fantasy.
---------------


My response :  Your comment/reaction is  INCORRECT.

Your correct comment/reaction would be  "Thanks for opening my eyes on the relevant facts I was not familiar with".


And had you comment in this way,  I'd tell you that this linear theory  DOESN'T  WORK  AT  ALL
for high-concentrated mixtures of salts in liquids.

            It works for low-concentrated salts mixtures in liquids,   O-N-L-Y.


Let me repeat it  one more time :   your reaction goes across your interests.


----------------

Alan,  on your comment
    "There's no reason to assume it's in NaCl in water"
I have good answer:
    But there is nothing in the problem that prohibits me to make this assumption . . . "


I always become very angry when see such  "problems".
Their creators do not know subject at all,  and all what they do is spreading  FAKES  in the Internet,  instead of spreading knowledge.



Answer by Alan3354(69443)

Actually, there's a lot a variation in the %age of salts in solution.
There's no reason to assume it's in NaCl in water.
=====================================
A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is 35% salt and Solution B is 85% salt. She wants to obtain 130 ounces of a mixture that is 80% salt. How many ounces of each solution should she use?
----------------------
A = amount of 35%
B = amount of 85%
----
A + B = 130 --- total solution
35A + 85B = 130*80 ---- total salt
---
Can you do the rest?
==========================
IDK if there is any salt that can be dissolved in any solvent at 85%.
I don't know why anyone would be angered by the question, tho.


Answer by greenestamps(13200)



We will overlook the scientific fact that you can't have an 80% salt solution (assuming you mean regular NaCl); we will simply treat this as a typical mixture problem.

Here is the fastest way to the answer to a mixture problem like this, without using any variation of the standard algebraic approach outlined by the other tutor.

(1) The 80% target is 9/10 of the way from 35% to 85%. (Picture the three percentages on a number line; 35 to 80 is a difference of 45; 35 to 85 is a difference of 50; 45/50 = 9/10.)
(2) Therefore, 9/10 of the mixture needs to be the higher percentage ingredient.

ANSWER: 9/10 of 130 ounces, or 117 ounces, of the 85% solution; the other 13 ounces is the 35% solution.

Answer by josgarithmetic(39617)

Make it a physical dry blend and only then the question makes sense.